The area of a rectangular banner should be 6.75m^2. If the length is to be thrice the width, what are its dimensions? (Show your solution)
Answer by mananth(16946)

let width be x
length = 3x
..
Area = L*W
area = x*3x
3x^2 = 6.75 m^2
/3
x^2= 6.75/3
x^2= 2.25
x= 1.5 m the width
...
Length = 3x
3* 1.5 = 4.5 m
Check
1.5*4.5= 6.75 sq.m